[Ncep.nhc.nco_contacts] TAFB Shift Log -- Technical Edition -- Wed May 10, 2017

TAFB Shift Log nhc.tsbadmin at noaa.gov
Wed May 10 22:44:35 UTC 2017


TAFB Shift Log for Wed May 10 2017 18:44:31 GMT-0400 (EDT)

Shift: Swing Shift

Pacific Forecaster: J Lewitsky

Atlantic Forecaster: M Nelson

Surface Analyst: N Ramos

Technical Issues Summary:

GFE/NCP issues:

N-AWIPS issues: Performance issues with STARFISH continue.

Other issues:

Miscellaneous issues: The cable boxes on the TAFB situational awareness  
monitor had to be unplugged and reset after they were turned off during  
b-roll shooting earlier in the day.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170510/06352db2/attachment-0001.html 


More information about the Ncep.nhc.nco_contacts mailing list